# Bifurcation mechanism of quasihalo orbit from Lissajous orbit

### v2

@@ 3.2 Solving for undetermined coefficients
When $(k,m)=(1,0)$, $x_{ijkm}$, $y_{ijkm}$, and $z_{ijkm}$ are coupled. Thus, $x_{ijkm}$ can be set zero, $z_{ijkm}=\eta x_{ijkm}=0$. In this case, (21) is turned to linear equations (24). Then, $y_{ijkm}$ and $\omega_{i-1j}$ are solved by the first two equations of (24). $d_{i-1j}$ is obtained from $\eta{d_{i-1j}}=-{\bar{r}_{ijkm}}-\eta\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$.
$-2{\omega_{00}}{y_{ijkm}}-2\left({{\omega_{00}}+\kappa}\right){\omega_{i-1j}}$ $={{\bar{p}}_{ijkm}}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$ (24)
$+({c_{2}}-1-\omega_{00}^{2}){y_{ijkm}}-2\left({\kappa{\omega_{00}}+1}\right){\omega_{i-1j}}$ $={{\bar{q}}_{ijkm}}+\kappa\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$
$-\eta{d_{i-1j}}$ $={{\bar{r}}_{ijkm}}{\rm{+}}\eta\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}.$
When $(k,m)=(0,1)$, the coefficient of $z_{ijkm}$ is zero, and therefore $z_{ijkm}$ is set to zero. In this case, (21) simplifies into linear equations (25). Then, $x_{ijkm}$ and $y_{ijkm}$ are solved by the first two equations of (25). $v_{ij-1}$ is obtained from $-2{v_{00}}{v_{ij-1}}={\bar{r}_{ijkm}}+{d_{00}}{x_{ijpq}}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}$.
$-(v_{00}^{2}+1+2{c_{2}}){x_{ijkm}}-2{v_{00}}{y_{ijkm}}$ $={{\bar{p}}_{ijkm}}$ (25)
$-2{v_{00}}{x_{ijkm}}+({c_{2}}-1-v_{00}^{2}){y_{ijkm}}$ $={{\bar{q}}_{ijkm}}$
$-{d_{00}}{x_{ijkm}}-2{v_{00}}{v_{ij-1}}$ $={{\bar{r}}_{ijkm}}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}.$
@@ 4 Results
In this section, the third-order analytical solution for central manifolds around the collinear libration points in the RTBP with arbitrary system parameter $\mu$ is derived. Moreover, the construction of the series solution up to a certain order $n$ is implemented for the given system parameter $\mu$, such as the Sun-Earth (S-E) system ($\mu=3.040423398444176\times 10^{-6}$) or Earth-Moon (E-M) system ($\mu=1.2150581918706896\times 10^{-2}$), utilizing the C++ 17 programming language.
@@ 4.1 Third-order analytical solution
$x$ $=\alpha\cos{\theta_{1}}+\left({{a_{21}}+{a_{22}}{\eta^{2}}}\right){\alpha^{2}}+\left({{a_{23}}+{a_{24}}{\eta^{2}}}\right){\alpha^{2}}\cos 2{\theta_{1}}$ (26)
$+{a_{25}}\eta\alpha\beta\cos({\theta_{1}}+{\theta_{2}})+{a_{26}}\eta\alpha\beta\cos({\theta_{1}}-{\theta_{2}})$
$+{a_{27}}{\beta^{2}}+{a_{28}}{\beta^{2}}\cos 2{\theta_{2}}$
$+({{a_{31}}{\eta^{4}}+{a_{32}}{\eta^{2}}+{a_{33}}}){\alpha^{3}}\cos 3{\theta_{1}}+({{a_{34}}{\eta^{3}}+{a_{35}}\eta}){\alpha^{2}}\beta\cos{\theta_{2}}$
$+({{a_{36}}{\eta^{3}}+{a_{37}}\eta}){\alpha^{2}}\beta\cos(2{\theta_{1}}+{\theta_{2}})$
$+({{a_{38}}{\eta^{3}}+{a_{39}}\eta}){\alpha^{2}}\beta\cos(2{\theta_{1}}-{\theta_{2}})$
$+({{a_{310}}{\eta^{2}}+{a_{311}}})\alpha{\beta^{2}}\cos({\theta_{1}}+2{\theta_{2}})$
$+({{a_{312}}{\eta^{2}}+{a_{313}}})\alpha{\beta^{2}}\cos({\theta_{1}}-2{\theta_{2}})$
$+{a_{314}}\eta{\beta^{3}}\cos{\theta_{2}}+{a_{315}}\eta{\beta^{3}}\cos 3{\theta_{2}},$
$y$ $=\kappa\alpha\sin{\theta_{1}}+({{b_{21}}+{b_{22}}{\eta^{2}}}){\alpha^{2}}\sin 2{\theta_{1}}+{b_{23}}\eta\alpha\beta\sin({\theta_{1}}+{\theta_{2}})$
$+{b_{24}}\eta\alpha\beta\sin({\theta_{1}}-{\theta_{2}})+{b_{25}}{\beta^{2}}\sin 2{\theta_{2}}$
$+({{b_{31}}{\eta^{4}}+{b_{32}}{\eta^{2}}+{b_{33}}}){\alpha^{3}}\sin{\theta_{1}}+({{b_{34}}{\eta^{4}}+{b_{35}}{\eta^{2}}+{b_{36}}}){\alpha^{3}}\sin 3{\theta_{1}}$
$+({{b_{37}}{\eta^{3}}+{b_{38}}\eta}){\alpha^{2}}\beta\sin{\theta_{2}}+({{b_{39}}{\eta^{3}}+{b_{310}}\eta}){\alpha^{2}}\beta sin(2{\theta_{1}}+{\theta_{2}})$
$+({{b_{311}}{\eta^{3}}+{b_{312}}\eta}){\alpha^{2}}\beta sin(2{\theta_{1}}-{\theta_{2}})+({{b_{313}}{\eta^{2}}+{b_{314}}})\alpha{\beta^{2}}\sin{\theta_{1}}$
$+({{b_{315}}{\eta^{2}}+{b_{316}}})\alpha{\beta^{2}}\sin({\theta_{1}}+2{\theta_{2}})$
$+({{b_{317}}{\eta^{2}}+{b_{318}}})\alpha{\beta^{2}}\sin({\theta_{1}}-2{\theta_{2}})$
$+{b_{319}}\eta{\beta^{3}}\sin{\theta_{2}}+{b_{320}}\eta{\beta^{3}}\sin 3{\theta_{2}},$
$z$ $=\eta\alpha\cos{\theta_{1}}+\left({{d_{21}}+{d_{22}}{\eta^{2}}}\right)\eta{\alpha^{2}}+\left({{d_{23}}+{d_{24}}{\eta^{2}}}\right)\eta{\alpha^{2}}\cos 2{\theta_{1}}$
$+({{d_{25}}+{d_{26}}{\eta^{2}}})\alpha\beta\cos({\theta_{1}}+{\theta_{2}})+({{d_{27}}+{d_{28}}{\eta^{2}}})\alpha\beta\cos({\theta_{1}}-{\theta_{2}})$
$+{d_{29}}\eta{\beta^{2}}+{d_{210}}\eta{\beta^{2}}\cos 2{\theta_{2}}+({{d_{31}}{\eta^{5}}+{d_{32}}{\eta^{3}}+{d_{33}}\eta}){\alpha^{3}}\cos 3{\theta_{1}}$
$+({{d_{34}}{\eta^{4}}+{d_{35}}{\eta^{2}}+{d_{36}}}){\alpha^{2}}\beta\cos(2{\theta_{1}}+{\theta_{2}})$
$+({{d_{37}}{\eta^{4}}+{d_{38}}{\eta^{2}}+{d_{39}}}){\alpha^{2}}\beta\cos(2{\theta_{1}}-{\theta_{2}})$
$+\left({{d_{310}}{\eta^{3}}+{d_{311}}\eta}\right)\alpha{\beta^{2}}\cos({\theta_{1}}+2{\theta_{2}})$
$+({{d_{312}}{\eta^{3}}+{d_{313}}\eta})\alpha{\beta^{2}}\cos({\theta_{1}}-2{\theta_{2}})$
$+({{d_{314}}{\eta^{2}}+{d_{315}}}){\beta^{3}}\cos 3{\theta_{2}},$
@@
$\omega$ $={\omega_{0}}+{\omega_{20}}+{\omega_{02}}={e_{31}}{\alpha^{2}}{\eta^{4}}+\left({{e_{32}}{\alpha^{2}}+{e_{33}}{\beta^{2}}}\right){\eta^{2}}+{e_{34}}{\alpha^{2}}+{e_{35}}{\beta^{2}}+{\omega_{0}}$ (27)
$v$ $={v_{0}}+{v_{20}}+{v_{02}}={e_{36}}{\alpha^{2}}{\eta^{4}}+\left({{e_{37}}{\alpha^{2}}+{e_{38}}{\beta^{2}}}\right){\eta^{2}}+{e_{39}}{\alpha^{2}}+{e_{310}}{\beta^{2}}+{v_{0}},$
@@
$\eta\Delta$ $=\eta\left({{d_{00}}+{d_{20}}{\alpha^{2}}+{d_{02}}{\beta^{2}}}\right)$ (28)
$=\eta\left[{{l_{1}}{\eta^{4}}{\alpha^{2}}+\left({{l_{2}}{\alpha^{2}}+{l_{3}}{\beta^{2}}}\right){\eta^{2}}+{l_{4}}{\alpha^{2}}+{l_{5}}{\beta^{2}}-\left({\omega_{0}^{2}-v_{0}^{2}}\right)}\right]=0$
where $a_{ij}$, $b_{ij}$, $d_{ij}$, $e_{ij}$, and $l_{i}$ are constant as provided in Appendix 5.1.
@@
It is not difficult to calculate and verify that $l_{1}>0,l_{2}<0,l_{3}<0,l_{4}>0$ and $l_{5}<0$ for the three collinear libration points with all system parameter $\mu\in\left({0,0.5}\right]$. Hence, (30) is a hyperbolic equation where a bifurcation occurs and suitable non-zero real solutions exist for (29) when $\Delta(\alpha,\beta)$ as illustrated in Fig. 1.
@@
The third-order solution (26) serves as an initial approximation to the center manifolds of around the collinear libration points in the RTBP. A higher-order series solution is required for real space missions. In such case, the bifurcation equation is no longer a hyperbola. Figure 1 just provides a basic outline of the feasible region for $\eta$, and in fact, its real feasible region is more intricate than depicted in Fig. 1. A higher-order numerical feasible region of $\eta$ will be presented in the following section.
Figure 2: Feasible region of $\eta$ for center manifolds of $L_{1}$ in the Sun-Earth system.
Figure 3: Distribution of $\eta$ within feasible region for center manifolds of $L_{1}$.

### v3

@@ 3.2 Solving for undetermined coefficients
When $(k,m)=(1,0)$, $x_{ijkm}$, $y_{ijkm}$, and $z_{ijkm}$ are couple. Thus, $x_{ijkm}$ can be set zero, $z_{ijkm}=\eta x_{ijkm}=0$. In this case, (21) is turned to linear equations (24). Then, $y_{ijkm}$ and $\omega{i-1j}$ are solved by the first two equations of (24). $d_{i-1j}$ is obtained from ${d_{i-1j}}=-{\bar{r}_{ijkm}}-\eta\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}$.
$\begin{array}[]{l}-2{\omega_{00}}{y_{ijkm}}-2\left({{\omega_{00}}+\kappa}\right){\omega_{i-1j}}={{\bar{p}}_{ijkm}}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}\\ +({c_{2}}-1-\omega_{00}^{2}){y_{ijkm}}-2\left({\kappa{\omega_{00}}+1}\right){\omega_{i-1j}}={{\bar{q}}_{ijkm}}+\kappa\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}\\ -{d_{i-1j}}={{\bar{r}}_{ijkm}}{\rm{+}}\eta\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{\omega_{{i_{1}}{j_{1}}}}{\omega_{{i_{2}}{j_{2}}}}}.\end{array}$ (24)
When $(k,m)=(0,1)$, the coefficient of $z_{ijkm}$ is zero, and therefore $z_{ijkm}$ is set zero. In this case, (21) simplifies into linear equations (25). Then, $x_{ijkm}$ and $y_{ijkm}$ are solved by the first two equations of (25). $v_{ij-1}$ is obtained from $-2{v_{00}}{v_{ij-1}}={\bar{r}_{ijkm}}+{d_{00}}{x_{ijpq}}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}$.
$\begin{array}[]{l}-(v_{00}^{2}+1+2{c_{2}}){x_{ijkm}}-2{v_{00}}{y_{ijkm}}={{\bar{p}}_{ijkm}}\\ -2{v_{00}}{x_{ijkm}}+({c_{2}}-1-v_{00}^{2}){y_{ijkm}}={{\bar{q}}_{ijkm}}\\ -{d_{00}}{x_{ijkm}}-2{v_{00}}{v_{ij-1}}={{\bar{r}}_{ijkm}}+\sum\limits_{{i_{1}},{j_{1}},{i_{2}},{j_{2}}}{C_{2}^{1-{\delta_{{i_{1}}{i_{2}}}}{\delta_{{j_{1}}{j_{2}}}}}{v_{{i_{1}}{j_{1}}}}{v_{{i_{2}}{j_{2}}}}}.\end{array}$ (25)
@@ 4 Results
In this section, the third-order analytical solution for central manifolds around the collinear libration points in the RTBP with arbitrary system parameter $\mu$ is derived. Moreover, the construction of the series solution up to a certain order $n$ is implemented for the given system parameter $\mu$, such as the Sun-Earth system ($\mu$ = 3.040423398444176e-6) or Earth-Moon system ($\mu$ = 1.215058191870689e-2), utilizing the C++ 17 programming language.
@@ 4.1 Third-order analytical solution
$\begin{array}[]{l}x=\alpha\cos{\theta_{1}}+\left({{a_{21}}+{a_{22}}{\eta^{2}}}\right){\alpha^{2}}+\left({{a_{23}}+{a_{24}}{\eta^{2}}}\right){\alpha^{2}}\cos 2{\theta_{1}}\\ +{a_{25}}\eta\alpha\beta\cos({\theta_{1}}+{\theta_{2}})+{a_{26}}\eta\alpha\beta\cos({\theta_{1}}-{\theta_{2}})\\ +{a_{27}}{\beta^{2}}+{a_{28}}{\beta^{2}}\cos 2{\theta_{2}}\\ +({{a_{31}}{\eta^{4}}+{a_{32}}{\eta^{2}}+{a_{33}}}){\alpha^{3}}\cos 3{\theta_{1}}+({{a_{34}}{\eta^{3}}+{a_{35}}\eta}){\alpha^{2}}\beta\cos{\theta_{2}}\\ +({{a_{36}}{\eta^{3}}+{a_{37}}\eta}){\alpha^{2}}\beta\cos(2{\theta_{1}}+{\theta_{2}})\\ +({{a_{38}}{\eta^{3}}+{a_{39}}\eta}){\alpha^{2}}\beta\cos(2{\theta_{1}}-{\theta_{2}})\\ +({{a_{310}}{\eta^{2}}+{a_{311}}})\alpha{\beta^{2}}\cos({\theta_{1}}+2{\theta_{2}})\\ +({{a_{312}}{\eta^{2}}+{a_{313}}})\alpha{\beta^{2}}\cos({\theta_{1}}-2{\theta_{2}})\\ +{a_{314}}\eta{\beta^{3}}\cos{\theta_{2}}+{a_{315}}\eta{\beta^{3}}\cos 3{\theta_{2}},\\ y=\kappa\alpha\sin{\theta_{1}}+({{b_{21}}+{b_{22}}{\eta^{2}}}){\alpha^{2}}\sin 2{\theta_{1}}+{b_{23}}\eta\alpha\beta\sin({\theta_{1}}+{\theta_{2}})\\ +{b_{24}}\eta\alpha\beta\sin({\theta_{1}}-{\theta_{2}})+{b_{25}}{\beta^{2}}\sin 2{\theta_{2}}\\ +({{b_{31}}{\eta^{4}}+{b_{32}}{\eta^{2}}+{b_{33}}}){\alpha^{3}}\sin{\theta_{1}}+({{b_{34}}{\eta^{4}}+{b_{35}}{\eta^{2}}+{b_{36}}}){\alpha^{3}}\sin 3{\theta_{1}}\\ +({{b_{37}}{\eta^{3}}+{b_{38}}\eta}){\alpha^{2}}\beta\sin{\theta_{2}}+({{b_{39}}{\eta^{3}}+{b_{310}}\eta}){\alpha^{2}}\beta sin(2{\theta_{1}}+{\theta_{2}})\\ +({{b_{311}}{\eta^{3}}+{b_{312}}\eta}){\alpha^{2}}\beta sin(2{\theta_{1}}-{\theta_{2}})+({{b_{313}}{\eta^{2}}+{b_{314}}})\alpha{\beta^{2}}\sin{\theta_{1}}\\ +({{b_{315}}{\eta^{2}}+{b_{316}}})\alpha{\beta^{2}}\sin({\theta_{1}}+2{\theta_{2}})\\ +({{b_{317}}{\eta^{2}}+{b_{318}}})\alpha{\beta^{2}}\sin({\theta_{1}}-2{\theta_{2}})\\ +{b_{319}}\eta{\beta^{3}}\sin{\theta_{2}}+{b_{320}}\eta{\beta^{3}}\sin 3{\theta_{2}},\end{array}$
$\begin{array}[]{l}z=\eta\alpha\cos{\theta_{1}}+\left({{d_{21}}+{d_{22}}{\eta^{2}}}\right)\eta{\alpha^{2}}+\left({{d_{23}}+{d_{24}}{\eta^{2}}}\right)\eta{\alpha^{2}}\cos 2{\theta_{1}}\\ +({{d_{25}}+{d_{26}}{\eta^{2}}})\alpha\beta\cos({\theta_{1}}+{\theta_{2}})+({{d_{27}}+{d_{28}}{\eta^{2}}})\alpha\beta\cos({\theta_{1}}-{\theta_{2}})\\ +{d_{29}}\eta{\beta^{2}}+{d_{210}}\eta{\beta^{2}}\cos 2{\theta_{2}}+({{d_{31}}{\eta^{5}}+{d_{32}}{\eta^{3}}+{d_{33}}\eta}){\alpha^{3}}\cos 3{\theta_{1}}\\ +({{d_{34}}{\eta^{4}}+{d_{35}}{\eta^{2}}+{d_{36}}}){\alpha^{2}}\beta\cos(2{\theta_{1}}+{\theta_{2}})\\ +({{d_{37}}{\eta^{4}}+{d_{38}}{\eta^{2}}+{d_{39}}}){\alpha^{2}}\beta\cos(2{\theta_{1}}-{\theta_{2}})\\ +\left({{d_{310}}{\eta^{3}}+{d_{311}}\eta}\right)\alpha{\beta^{2}}\cos({\theta_{1}}+2{\theta_{2}})\\ +({{d_{312}}{\eta^{3}}+{d_{313}}\eta})\alpha{\beta^{2}}\cos({\theta_{1}}-2{\theta_{2}})\\ +({{d_{314}}{\eta^{2}}+{d_{315}}}){\beta^{3}}\cos 3{\theta_{2}},\end{array}$ (26)
@@
$\begin{array}[]{l}\omega={\omega_{0}}+{\omega_{20}}+{\omega_{02}}\\ ={e_{31}}{\alpha^{2}}{\eta^{4}}+\left({{e_{32}}{\alpha^{2}}+{e_{33}}{\beta^{2}}}\right){\eta^{2}}+{e_{34}}{\alpha^{2}}+{e_{35}}{\beta^{2}}+{\omega_{0}}\\ v={v_{0}}+{v_{20}}+{v_{02}}\\ ={e_{36}}{\alpha^{2}}{\eta^{4}}+\left({{e_{37}}{\alpha^{2}}+{e_{38}}{\beta^{2}}}\right){\eta^{2}}+{e_{39}}{\alpha^{2}}+{e_{310}}{\beta^{2}}+{v_{0}},\end{array}$ (27)
@@
$\begin{array}[]{l}\eta\Delta=\eta\left({{d_{00}}+{d_{20}}{\alpha^{2}}+{d_{02}}{\beta^{2}}}\right)\\ =\eta\left[{{l_{1}}{\eta^{4}}{\alpha^{2}}+\left({{l_{2}}{\alpha^{2}}+{l_{3}}{\beta^{2}}}\right){\eta^{2}}+{l_{4}}{\alpha^{2}}+{l_{5}}{\beta^{2}}-\left({\omega_{0}^{2}-v_{0}^{2}}\right)}\right]=0\end{array}$ (28)
where $a_{ij}$, $b_{ij}$, $d_{ij}$, $e_{ij}$, and $l_{i}$ are constant as provided in A.
@@
It is not difficult to calculate and verify that $l_{1}>0,l_{2}<0,l_{3}<0,l_{4}>0$ and $l_{5}<0$ for the three collinear libration points with all system parameter $\mu\in\left({0,0.5}\right]$. Hence, (30) is a hyperbolic equation where bifurcation occurs and suitable non-zero real solutions exist for (29) when $\Delta(\alpha,\beta)$ as illustrated in Fig. 1.
@@
Remark 4. The third-order solution (26) serves as an initial approximation to the center manifolds of around the collinear libration points in the RTBP. A higher-order series solution is required for real space missions. In such case, the bifurcation equation is no longer a hyperbola. Figure 1 just provides a basic outline of the feasible region for $\eta$, and in fact, its real feasible region is more intricate than depicted in Fig. 1. A higher-order numerical feasible region of $\eta$ will be presented in the following section.
Figure 2: Feasible region of coupling coefficient $\eta$ for center manifolds of $L_{1}$ in the Earth-Sun system with different orders $n$ = 3, 5, 7, 11, 15, 19. (In this case, $\mu$ = 3.040423398444176e-6, $\gamma_{1}$ = 1.00109772277814e-2) with $\alpha$ and $\beta$.
Figure 3: Contour maps of coupling coefficient $\eta$ within feasible region for center manifolds of $L_{1}$ in the Earth-Sun.